# LiveResearchBench: A Live Benchmark for User-Centric Deep Research in the Wild

### v1

@@ 1 Introduction
Deep research is the process of addressing complex, open-ended questions that require long-horizon, multi-step reasoning and planning. It involves exploring hundreds of live web sources and synthesizing them into comprehensive, citation-grounded, and structured outputs such as reports (OpenAI, 2025). This process closely mirrors how human researchers work—iteratively seeking, validating, and integrating evidence. Building systems capable of this marks an important frontier in AI: moving LLMs from simple chatbots to independent problem-solvers. However, progress is bottlenecked by a benchmarking and evaluation crisis: Which tasks are truly representative? How should ambiguities in query formulation be resolved? How can we reliably measure comprehensiveness (breadth and depth), consistency, factuality, and citation accuracy in long-form outputs while safeguarding against data contamination (Nguyen et al., 2025)?
Evaluating agentic tasks is challenging in general, and deep research is particularly difficult to benchmark due to its open-ended nature. To address this, we first propose four key principles for benchmarking deep research systems: tasks should be (a) user-centric, reflecting realistic information needs; (b) dynamic, requiring up-to-date information beyond parametric knowledge; (c) unambiguous, ensuring consistent interpretation across users; and (d) multi-faceted and search-intensive, requiring multi-hop search across diverse web sources and in-depth analysis. Recent benchmarks have introduced search-intensive tasks (Gou et al., 2025), but these are often static (and thus prone to contamination) (Bosse et al., 2025), domain-specific (Xu et al., 2025b; Patel et al., 2025), coarse-grained, or ambiguous (Du et al., 2025)—frequently omitting details such as the intended audience, required output format, or scope. Consequently, we still lack reliable tools to assess whether models are producing genuinely high-quality research or merely assembling plausible-sounding yet insufficient responses.
@@ 2 Related Work
Single- and Multi-Agent Deep Research Systems. Equipping LLMs with tools—often referred to as agents—for grounded answers to complex queries has become a major research direction. Pioneering systems such as Deep Research using o3 (OpenAI, 2025) combined browsing and code execution and generate long reports. Many successors have since emerged (Xu & Peng, 2025; Java et al., 2025b). Architectures generally fall into two broad categories. Single-agent systems place all tool-use decisions in one model. Examples include function-calling LLMs (Yang et al., 2025; Agarwal et al., 2025) and ReAct-style agents (Yao et al., 2022). Multi-agent systems (MAS) instead coordinate specialized roles such as planner, researcher, and writer in a predefined workflow. Proprietary multi-agent systems such as Grok 4 Heavy (xAI, 2025a) and Manus (Manus, 2025) remain undisclosed, while examples of open-source systems include Open Deep Research (Alzubi et al., 2025), DeerFlow (ByteDance, 2025), and OpenManus (Liang et al., 2025). Despite rapid progress, it is still unclear which dimensions favor MAS vs. single-agent systems, what fundamental bottlenecks remain, and whether MAS truly deliver consistent gains (Kapoor et al., 2025). The lack of standardized benchmarks and evaluation frameworks further prevents rigorous comparison. Our work directly addresses this gap by developing a comprehensive evaluation framework, enabling rigorous comparison of single- and multi-agent systems for dynamic and multi-faceted research tasks.
@@
Deep Research Benchmark. Several benchmarks have been proposed for evaluating deep research. DeepScholarBench (Patel et al., 2025) targets related-work generation; DeepResearch Bench (Du et al., 2025) include 100 short open-ended questions; Deep Research Bench (Bosse et al., 2025), LiveDRBench (Java et al., 2025a), and Mind2Web2 (Gou et al., 2025) focus on information-seeking tasks. While valuable, these benchmarks are typically domain-specific, ambiguous in query formulation, limited to short-form answers with low reasoning load, and static, making fair comparison difficult and preventing evaluation on evolving information. A comparison is shown in Figure 1. Recent efforts such as DeepResearchGym (Coelho et al., 2025) adopts short and static tasks, while ResearcherBench (Xu et al., 2025b) focuses narrowly on AI domain. In contrast, our LiveResearchBench provides user-centric, unambiguous, and time-varying tasks that are multi-faceted, search-intensive, and span diverse domains.
@@
Evaluation of Long-form Answer. Recent work has moved beyond simple reference-based metrics. Some benchmarks emphasize coverage and factual grounding (Laban et al., 2023; 2024; Huang et al., 2024; Xu et al., 2025a), such as PROXYQA (Tan et al., 2024) (sub-question coverage) and LongFact (Wei et al., 2024) (claim-level verification). Others target structure and overall quality, including LongEval (Wu et al., 2025) (plan-based vs. direct generation) and HelloBench (Que et al., 2024) (human-aligned protocols). However, these efforts typically isolate a single dimension. In contrast, DeepEval offers a comprehensive suite: report-level metrics (presentation, consistency) and fine-grained dimensions (depth, checklist-based coverage, citation association, rubric-tree accuracy). Crucially, DeepEval is reliable and human-aligned: instead of holistic scores, we use structured checklists (e.g., factual errors as atomic tests) and aggregate results, yielding stable grades that closely track expert judgments.
@@ 3.2 Benchmark Overview
Compared to prior deep research benchmarks, LiveResearchBench tasks are designed to be more unambiguous and realistic (Figure 1). For example, Deep Research Bench (Bosse et al., 2025) includes search-intensive queries such as How many IM and GM account closures did chess.com report for 2024?, but these are static, time-bounded, and require minimal reasoning—limiting their ability to evaluate evolving real-world research. Likewise, DeepResearch Bench (Du et al., 2025) often omits crucial elements such as the target audience, content and format requirements, or a clearly defined scope. Such omissions leave queries open to multiple interpretations, even across runs of the same model, undermining evaluation reliability.
@@ 4 DeepEval: A Comprehensive Evaluation Suite for Deep Research
As the metrics differ in granularity, no single evaluation protocol is sufficient. For metrics ❶-❺222❻Citation Accuracy requires verifying (statement, URL) pairs throughout the report. This task is naturally handled using the rubric tree protocol detailed in Section 4.2., we consider three options (Gu et al., 2025): (1) Checklist-based, where the judge evaluates the generated report against a checklist and provides a binary score (0/1) for each item in the checklist. (2) Pointwise (additive), where the judge aims to find concrete errors in the report according to the criteria. The final score is determined based on the number of errors. (3) Pairwise comparison, where the judge compares two reports side by side and selects the preferred one (or declares a tie). Based on preliminary experiments, we narrowed down the choice of protocol for each metric, and subsequently conducted a human alignment study on the selected protocols. As a result, we adopt Checklist-based for ❶ Presentation & Organization and ❸ Coverage & Comprehensiveness; Pointwise (additive) for ❷ Factual & Logical Consistency and ❺ Citation Association; Pairwise comparison for ❹ 0pt. These protocols achieve high agreement with human judgments (see Appendix B).
@@ 4.2 Fine-grained: Content-level Metrics
❹ Analysis Depth. Beyond coverage, a strong research report should provide substantive insights rather than superficial information gathering. We assess depth along five dimensions: (1) granularity of reasoning (causal chains), (2) multi-layer insights (trade-offs and implications), (3) critical evaluation (critiques and limitations), (4) analytical use of evidence, and (5) insight density. Ensemble judges conduct pairwise comparisons between the evaluated and baseline reports, assigning 1–5 scores to each along every dimension. To mitigate positional bias, we adopt the position-swap averaging method of Wang et al. (2023), scoring each pair twice with reversed order and averaging results. Dimension scores are summed, averaged across judges, and used to derive win rates: A win, B win, or Tie (if scores differ by ). Ties are excluded from the denominator. The human agreement rate is 92.5% with the judge ensemble. Detailed prompts can be seen in Appendix A.4.
@@ 5 Main Results and analysis
Agentic Systems. We conduct a comprehensive evaluation of 17 state-of-the-art open-sourced and proprietary agentic systems, which can usually be grouped into three categories: (1) Single-agent systems with web search capabilities, including GPT-5 (OpenAI, 2025a) , GPT-4.1 (OpenAI, 2024), GPT-5-mini (OpenAI, 2025b), Gemini 2.5 Pro (DeepMind, 2025b), Gemini 2.5 Flash (DeepMind, 2025a), Claude 4 Sonnet (Anthropic, 2025a), Claude 4.1 Opus (Anthropic, 2025b), Perplexity Sonar Reasoning (Perplexity, 2025a), and Perplexity Sonar Reasoning Pro (Perplexity, 2025b); (2) Single-agent deep research systems, which feature extended reasoning depth and longer thinking time, including OpenAI o3 Deep Research (OpenAI, 2025c), OpenAI o4-mini Deep Research (OpenAI, 2025d), Perplexity Sonar Deep Research (AI, 2025b), Grok-4 Deep Research (Expert) (xAI, 2025b), and Gemini Deep Research (DeepMind, 2025c); (3) Multi-agent deep research systems, which coordinate a team of specialized agents to decompose complex queries. These systems typically employ a planner to dispatch sub-agents for browsing, searching, synthesis, and verification, while maintaining a large shared context across tasks. This category includes Manus (Manus, 2025), Grok-4 Heavy Deep Research (xAI, 2025a), Open Deep Research (AI, 2025a), and Deerflow+.
Deerflow+ is our enhanced implementation of Deerflow (ByteDance, 2025) that resolves frequent failure modes and early termination by adding inline citations, reference mapping, and robust long-context management. Compared to Deerflow, Deerflow+ eliminates token-limit failures, retains more retrieved information, and exhibits stronger evidence structuring and citation discipline. Implementation details are provided in Appendix C; Figure 13 shows side-by-side outputs for the same prompt in Figure 1.

### v3

@@ Abstract
livedeepresearch.github.io
huggingface.co/datasets/Salesforce/LiveResearchBench
github.com/SalesforceAIResearch/LiveResearchBench
@@ 1 Introduction
Deep research is the process of addressing complex, open-ended questions that require long-horizon, multi-step reasoning and planning. It involves exploring hundreds of live web sources and synthesizing them into comprehensive, citation-grounded, and structured outputs such as reports (openai2025deepresearch). This process closely mirrors how human researchers work—iteratively seeking, validating, and integrating evidence. Building systems capable of this marks an important frontier in AI: moving LLMs from simple chatbots to independent problem-solvers. However, progress is bottlenecked by a benchmarking and evaluation crisis: Which tasks are truly representative? How should ambiguities in query formulation be resolved? How can we reliably measure comprehensiveness (breadth and depth), consistency, factuality, and citation accuracy in long-form outputs while safeguarding against data contamination (nguyen2025sfrdeepresearch)?
Evaluating agentic tasks is challenging in general, and deep research is particularly difficult to benchmark due to its open-ended nature. To address this, we first propose four key principles for benchmarking deep research systems: tasks should be (a) user-centric, reflecting realistic information needs; (b) dynamic, requiring up-to-date information beyond parametric knowledge; (c) unambiguous, ensuring consistent interpretation across users; and (d) multi-faceted and search-intensive, requiring multi-hop search across diverse web sources and in-depth analysis. Recent benchmarks have introduced search-intensive tasks (gou2025mindweb), but these are often static (and thus prone to contamination) (bosse2025deep), domain-specific (xu2025researcherbench; patel2025deepscholarbenchlivebenchmarkautomated), coarse-grained, or ambiguous (du2025deepresearch)—frequently omitting details such as the intended audience, required output format, or scope. Consequently, we still lack reliable tools to assess whether models are producing genuinely high-quality research or merely assembling plausible-sounding yet insufficient responses.
@@ 2 Related Work
Single- and Multi-Agent Deep Research Systems. Equipping LLMs with tools—often referred to as agents—for grounded answers to complex queries has become a major research direction. Pioneering systems such as Deep Research using o3 (openai2025deepresearch) combined browsing and code execution and generate long reports. Many successors have since emerged (deepresearch_survey_xu2025comprehensive; java2025characterizing_dr). Architectures generally fall into two broad categories. Single-agent systems place all tool-use decisions in one model. Examples include function-calling LLMs (yang2025qwen3; agarwal2025gpt) and ReAct-style agents (react_yao2022react). Multi-agent systems (MAS) instead coordinate specialized roles such as planner, researcher, and writer in a predefined workflow. Proprietary multi-agent systems such as Grok 4 Heavy (Grok4Heavy2025) and Manus (manus2025) remain undisclosed, while examples of open-source systems include Open Deep Research (alzubi2025opendeepsearch), DeerFlow (DeerFlow2025), and OpenManus (OpenManus2025). Despite rapid progress, it is still unclear which dimensions favor MAS vs. single-agent systems, what fundamental bottlenecks remain, and whether MAS truly deliver consistent gains (kapoor2025ai). The lack of standardized benchmarks and evaluation frameworks further prevents rigorous comparison. Our work directly addresses this gap by developing a comprehensive evaluation framework, enabling rigorous comparison of single- and multi-agent systems for dynamic and multi-faceted research tasks.
@@
Deep Research Benchmark. Several benchmarks have been proposed for evaluating deep research. DeepScholarBench (patel2025deepscholarbenchlivebenchmarkautomated) targets related-work generation; DeepResearch Bench (du2025deepresearch) include 100 short open-ended questions; Deep Research Bench (bosse2025deep), LiveDRBench (java2025characterizing), and Mind2Web2 (gou2025mindweb) focus on information-seeking tasks. While valuable, these benchmarks are typically domain-specific, ambiguous in query formulation, limited to short-form answers with low reasoning load, and static, making fair comparison difficult and preventing evaluation on evolving information. A comparison is shown in Figure 1. Recent efforts such as DeepResearchGym (coelho2025deepresearchgym) adopts short and static tasks, while ResearcherBench (xu2025researcherbench) focuses narrowly on AI domain. In contrast, our LiveResearchBench provides user-centric, unambiguous, and time-varying tasks that are multi-faceted, search-intensive, and span diverse domains.
@@
Evaluation of Long-form Answer. Recent work has moved beyond simple reference-based metrics. Some benchmarks emphasize coverage and factual grounding (laban-etal-2023-summedits; laban2024SummHay; huang-etal-2024-embrace; xu-etal-2025-context), such as PROXYQA (tan2024proxyqaalternativeframeworkevaluating) (sub-question coverage) and LongFact (wei2024longformfactualitylargelanguage) (claim-level verification). Others target structure and overall quality, including LongEval (wu2025longevalcomprehensiveanalysislongtext) (plan-based vs. direct generation) and HelloBench (que2024hellobenchevaluatinglongtext) (human-aligned protocols). However, these efforts typically isolate a single dimension. In contrast, DeepEval offers a comprehensive suite: report-level metrics (presentation, consistency) and fine-grained dimensions (depth, checklist-based coverage, citation association, rubric-tree accuracy). Crucially, DeepEval is reliable and human-aligned: instead of holistic scores, we use structured checklists (e.g., factual errors as atomic tests) and aggregate results, yielding stable grades that closely track expert judgments.
@@ 3.2 Benchmark Overview
Compared to prior deep research benchmarks, LiveResearchBench tasks are designed to be more unambiguous and realistic (Figure 1). For example, Deep Research Bench (bosse2025deep) includes search-intensive queries such as How many IM and GM account closures did chess.com report for 2024?, but these are static, time-bounded, and require minimal reasoning—limiting their ability to evaluate evolving real-world research. Likewise, DeepResearch Bench (du2025deepresearch) often omits crucial elements such as the target audience, content and format requirements, or a clearly defined scope. Such omissions leave queries open to multiple interpretations, even across runs of the same model, undermining evaluation reliability.
@@ 4 DeepEval: A Comprehensive Evaluation Suite for Deep Research
As the metrics differ in granularity, no single evaluation protocol is sufficient. For metrics ❶-❺222❻Citation Accuracy requires verifying (statement, URL) pairs throughout the report. This task is naturally handled using the rubric tree protocol detailed in Section 4.2., we consider three options (gu2025surveyllmasajudge): (1) Checklist-based, where the judge evaluates the generated report against a checklist and provides a binary score (0/1) for each item in the checklist. (2) Pointwise (additive), where the judge aims to find concrete errors in the report according to the criteria. The final score is determined based on the number of errors. (3) Pairwise comparison, where the judge compares two reports side by side and selects the preferred one (or declares a tie). Based on preliminary experiments, we narrowed down the choice of protocol for each metric, and subsequently conducted a human alignment study on the selected protocols. As a result, we adopt Checklist-based for ❶ Presentation & Organization and ❸ Coverage & Comprehensiveness; Pointwise (additive) for ❷ Factual & Logical Consistency and ❺ Citation Association; Pairwise comparison for ❹ 0pt. These protocols achieve high agreement with human judgments (see Appendix B).
@@ 4.2 Fine-grained: Content-level Metrics
❹ Analysis Depth. Beyond coverage, a strong research report should provide substantive insights rather than superficial information gathering. We assess depth along five dimensions: (1) granularity of reasoning (causal chains), (2) multi-layer insights (trade-offs and implications), (3) critical evaluation (critiques and limitations), (4) analytical use of evidence, and (5) insight density. Ensemble judges conduct pairwise comparisons between the evaluated and baseline reports, assigning 1–5 scores to each along every dimension. To mitigate positional bias, we adopt the position-swap averaging method of wang2023largefair, scoring each pair twice with reversed order and averaging results. Dimension scores are summed, averaged across judges, and used to derive win rates: A win, B win, or Tie (if scores differ by ). Ties are excluded from the denominator. The human agreement rate is 92.5% with the judge ensemble. Detailed prompts can be seen in Appendix A.4.
@@ 5 Main Results and analysis
Agentic Systems. We conduct a comprehensive evaluation of 17 state-of-the-art open-sourced and proprietary agentic systems, which can usually be grouped into three categories: (1) Single-agent systems with web search capabilities, including GPT-5 (openai2025gpt5) , GPT-4.1 (openai2024gpt41), GPT-5-mini (openai2025gpt5mini), Gemini 2.5 Pro (deepmind2025gemini25pro), Gemini 2.5 Flash (deepmind2025gemini25flash), Claude 4 Sonnet (anthropic2025claudesonnet), Claude 4.1 Opus (anthropic2025opus41), Perplexity Sonar Reasoning (perplexity2025sonar), and Perplexity Sonar Reasoning Pro (perplexity2025sonarpro); (2) Single-agent deep research systems, which feature extended reasoning depth and longer thinking time, including OpenAI o3 Deep Research (openai2025o3deep), OpenAI o4-mini Deep Research (openai2025o4minideep), Perplexity Sonar Deep Research (perplexity2025sonardeep), Grok-4 Deep Research (Expert) (xai2025grok4expert), and Gemini Deep Research (deepmind2025geminideep); (3) Multi-agent deep research systems, which coordinate a team of specialized agents to decompose complex queries. These systems typically employ a planner to dispatch sub-agents for browsing, searching, synthesis, and verification, while maintaining a large shared context across tasks. This category includes Manus (manus2025), Grok-4 Heavy Deep Research (Grok4Heavy2025), Open Deep Research (langchain2025opendeep), and Deerflow+.
Deerflow+ is our enhanced implementation of Deerflow (DeerFlow2025) that resolves frequent failure modes and early termination by adding inline citations, reference mapping, and robust long-context management. Compared to Deerflow, Deerflow+ eliminates token-limit failures, retains more retrieved information, and exhibits stronger evidence structuring and citation discipline. Implementation details are provided in Appendix C; Figure 13 shows side-by-side outputs for the same prompt in Figure 1.